# DeblurSDI: Blind Image Deblurring Using Self-diffusion

### v1

@@ 4.1 Network architecture and training
Due to the low-dimensional nature of the blur kernel, we employ a fully-connected network (FCN) to implement the kernel generator, $G_{\phi}$. To ensure the output corresponds to a physical blur kernel, a softmax activation is applied to the final layer, enforcing non-negativity and a sum-to-one constraint. The 1D output of $G_{\phi}$ is subsequently reshaped into a 2D blur kernel. Besides, we introduced standard mode for ablation study where the latent vector $z$ is sampled from a normal distribution and kept fixed during training, and diffusion mode where the $\mathbf{z}_{t}$ evolves through the self-diffusion process. $K$ is the kernel size, $n$ is the number of hidden layers and $H_{d}$ is the hidden dimension size. Figure 1 shows how the architecture of $G_{\phi}$ affects the performance of DeblurSDI.
@@
For the image denoiser $D_{\theta}$, we employ an encoder-decoder network with skip connections, following a U-Net-like structure. The network consists of five hierarchical levels. Each level in the encoder path consists of two convolutional blocks and a stride-2 convolution for downsampling. Correspondingly, the decoder path uses bilinear upsampling. Skip connections concatenate features from each encoder level to the corresponding decoder level. Non-Local Blocks are integrated into the deeper encoder levels (levels 3, 4, and 5) to capture long-range dependencies. The architecture is detailed in Appendix B.
Figure 1: Ablaion study on the network for blur kernel estimation. This figure compares the performance of the “Standard” mode against the “Diffusion” mode architecture with a varying number of hidden layers ($n=1$ to $n=5$). The results clearly indicate that increasing the depth of the kernel generator network leads to a more accurate kernel estimation
4.2 Hyperparameters sensitivity
The number of outer diffusion steps $T$ and inner optimization iterations $S$ are two critical hyperparameters that directly impact both the reconstruction quality and computational cost. Intuitively, more diffusion steps allow for a finer coarse-to-fine reconstruction, while more inner iterations enable better convergence of the networks at each step. However, increasing either parameter also leads to longer runtimes. To evaluate the sensitivity of our method to these parameters, we conduct experiments varying $T$ from 10 to 40 and $S$ from 25 to 500. As shown in Figure 2, performance improves with higher values of $T$ and $S$. The most significant gains occur when increasing $T$ from 10 to 30. However, the improvements tend to saturate beyond certain thresholds (e.g., $T=30$, $S=400$), with the performance curve for $T=40$ closely tracking that of $T=30$. This indicates that our approach can achieve strong deblurring performance without requiring excessively high iteration counts.
Figure 2: Sensitivity of noise steps, $T$, and inner iterations, $S$. The graphs show the SSIM (left) and PSNR (right) scores for different numbers of outer diffusion steps ($T\in\{10,20,30,40\}$) and inner optimization iterations ($S\in\{25,...,500\}$).
4.3 Deblurring process
4.4 Robustness to kernel size
The joint estimation of the image denoiser and the blur kernel often collapses to trivial solutions, such as Dirac kernels or reproducing the blurred image itself, especially when the chosen kernel size is incompatible with the image content. Larger kernels are harder to be recovered accurately, while smaller kernels may fail to capture long-range motion. For this reason, SelfDeblur (Ren et al., 2020) carefully selects the kernel size for each image. In contrast, our method exhibits much greater robustness. As shown in Figure 4, we evaluate ten different kernel sizes from 15 to 33, and compare the performance of several approaches. Our method not only achieves consistently superior performance across all kernel sizes but also demonstrates remarkable stability.
5 Evaluation
5.1 Dataset
To systematically evaluate the performance of our method, we collected four datasets including Levin (Levin et al., 2007), Cho (Cho & Lee, 2009), Kohler (Köhler et al., 2012), and FFHQ (Karras et al., 2019). The first three datasets are widely used benchmarks for blind image deblurring, containing various synthetic blur kernels applied to natural images. The FFHQ dataset consists of 20 random selected human face images and 4 blur kernels from Cho (Cho & Lee, 2009). For our experiments, we use $T=30$ outer reverse diffusion steps. At each step, we run $S=200$ inner optimization iterations. For each dataset, we use the provided blur kernels and generate blurred images by convolving the sharp images with these kernels to simulate real-world conditions. Each blur kernel is applied to every image in the dataset. The details of our comprehensive datasets are shown in Table 2.
Table 2: Details of datasets for evaluation
Image size Kernel Size Pairs
Levin1 $\times255$ (19, 17, 15, 27, 13, 21, 23, 23) 32
[…]
Kohler3 $\times800$ (16, 14, 9, 13, 29, 17, 19, 98, 102, 62, 40, 29) 36
FFHQ4 $\times256$ (27, 23, 19, 21) 80
Total - - 128
1(Levin et al., 2007) 2(Cho & Lee, 2009) 3(Köhler et al., 2012) 4(Karras et al., 2019)
5.2 Comparison with Other Methods
We compare our DeblurSDI method with several other blind deblurring approaches, including Phase-Only (Pan et al., 2019), FFT-ReLU Deblur (Al Radi et al., 2025), SelfDeblur (Ren et al., 2020), and FastDiffusionEM (Laroche et al., 2024). We evaluate the performance of each method using Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM) on the four datasets mentioned above. The quantitative results are summarized in Table 3. Rather than setting specific kernel size for each image (Ren et al., 2020), we set a fixed kernel size for each dataset, i.e., 27 for Levin, 33 for Cho, Kohler and FFHQ. As shown in the table, our DeblurSDI consistently outperforms all compared methods across all datasets, achieving significant improvements in both PSNR and SSIM metrics. This demonstrates the effectiveness and generalizability of our self-diffusion approach in recovering sharp images and accurate blur kernels.
Table 3: Quantitative results of blind deblurring performance (PSNR/SSIM) on four datasets.
Phase-Only1 FFT-ReLU Deblur2 SelfDeblur3 FastDiffusionEM4 DeblurSDI (Ours)
Levina 20.68/0.6061 15.56/0.3845 25.06/0.7301 16.55/0.4005 31.85/0.7911
Chob 19.89/0.6746 18.73/0.6546 20.37/0.6844 15.39/0.4687 28.73/0.8859
Kohlerc 28.23/0.8092 25.33/0.7140 21.97/0.5995 18.85/0.4813 29.17/0.7653
FFHQd 25.80/0.7904 21.71/0.6579 19.82/0.5563 15.59/0.3592 33.90/0.9064
1 (Pan et al., 2019), 2 (Al Radi et al., 2025), 3 (Ren et al., 2020), 4 (Laroche et al., 2024)
a (Levin et al., 2007), b (Cho & Lee, 2009), c (Köhler et al., 2012), d (Karras et al., 2019)
Figure 5: Deblurring results on the FFHQ dataset (Karras et al., 2019).
Figure 5 shows deblurring results of different methods on the FFHQ dataset (Karras et al., 2019). For each estimated image, the recovered kernel is displayed at the top-left corner, and three zoomed-in regions highlight fine details. As observed, FastDiffusionEM (Laroche et al., 2024) performs the worst. Despite being pre-trained on the FFHQ dataset, its performance remains poor: the estimated kernels degenerate into trivial point- or line-like structures. Without reliable kernel estimation, even a strong image prior cannot yield satisfactory deblurring. In contrast, other four optimization-based methods produce visibly superior results. PhaseOnly (Pan et al., 2019) and FFTReLU Deblur (Al Radi et al., 2025) provide good results in some cases, but lack of robustness when facing various blur kernels. SelfDeblur (Ren et al., 2020) shows the most promising performance out of previous methods, yet still suffers from poor generalizability, reconstruction shiftting and color distortion. The proposed self-diffusion-based DeblurSDI method consistently outperforms all compared methods, achieving significant improvements in the long-standing shiftting and robustness issues of blind deblurring techniques. Moreover, DeblurSDI is also able to recover accurate blur kernels, which gurantees the performance of the method. More results of other three datasets can be found in Appendix C.
6 Conclusion
This paper presents a novel self-diffusion-based approach for blind image deblurring, which we call DeblurSDI. Our method leverages the self-diffusion principle to recover accurate blur kernels and sharp images in a single framework. Experimental results on four benchmark datasets show that DeblurSDI consistently outperforms other blind deblurring methods on various datasets, demonstrating its effectiveness and generalizability in recovering sharp images and accurate blur kernels.
Appendix A LLM Usage Statement
We used a large language models only to refine grammar and improve the clarity of language in this manuscript. No part of the research ideation, experiment design, or analysis was performed by an LLM.
Appendix B Architecture of the Image Denoiser ($D_{\theta}$)
@@
Appendix C More Results
Figure 6: Deblurring results on the Cho dataset (Cho & Lee, 2009).
Figure 7: Deblurring results on the Kohler dataset (Köhler et al., 2012).
Figure 8: Deblurring results on the Levin dataset (Levin et al., 2007).

### v2

@@ 4.2 Network architecture and training
Due to the low-dimensional nature of the blur kernel, we employ a fully-connected network (FCN) to implement the kernel generator, $G_{\phi}$. To ensure the output corresponds to a physical blur kernel, a softmax activation is applied to the final layer, enforcing non-negativity and a sum-to-one constraint. The 1D output of $G_{\phi}$ is subsequently reshaped into a 2D blur kernel. Besides, we introduced standard mode for ablation study where the latent vector $z$ is sampled from a normal distribution and kept fixed during training, and diffusion mode where the $\mathbf{z}_{t}$ evolves through the self-diffusion process. $K$ is the kernel size, $n$ is the number of hidden layers and $H_{d}$ is the hidden dimension size, Table 2 shows the architecture in details.
@@
For the image denoiser $D_{\theta}$, we employ an encoder-decoder network with skip connections, following a U-Net-like structure. The network consists of five hierarchical levels. Each level in the encoder path consists of two convolutional blocks and a stride-2 convolution for downsampling. Correspondingly, the decoder path uses bilinear upsampling. Skip connections concatenate features from each encoder level to the corresponding decoder level. Non-Local Blocks are integrated into the deeper encoder levels (levels 3, 4, and 5) to capture long-range dependencies. The architecture is detailed as in Table 3.
@@
4.3 Experiments
In the following, we present the deblurring process described in Algorithm 1 and evaluated DeblurSDI from different perspectives, including robustness to the choice of PSF generator, sensitivity to hyperparameters, and noise degradation.
Figure 2: Sensitivity of noise scheduled steps, $T$, and inner iterations, $S$. The graphs show the SSIM (left) and PSNR (right) scores for different numbers of outer diffusion steps ($T\in\{10,20,30,40\}$) and inner optimization iterations ($S\in\{25,...,500\}$).
Hyperparameters sensitivity:
The number of outer diffusion steps $T$ and inner optimization iterations $S$ in Algorithm 1 are two critical hyperparameters that directly impact both the reconstruction quality and computational cost. Intuitively, more diffusion steps allow for a finer coarse-to-fine reconstruction, while more inner iterations enable better convergence of the networks at each step. However, increasing either parameter also leads to longer runtimes. To evaluate the sensitivity of our method to these parameters, we conduct experiments varying $T$ from 10 to 40 and $S$ from 25 to 500. As shown in Figure 2, performance improves with higher values of $T$ and $S$. The most significant gains occur when increasing $T$ from 10 to 30. However, the improvements tend to saturate beyond certain thresholds (e.g., $T=30$, $S=400$), with the performance curve for $T=40$ closely tracking that of $T=30$. This indicates that our approach can achieve strong deblurring performance without requiring excessively high iteration counts.
Recovering process:
Architecture of kernel generator:
Figure 4 shows how the architecture of PSF kernel generator $G_{\phi}$ affects the performance of DeblurSDI. The “Standard” mode is $G_{\phi}$ without noise schedule, while “k-diff” stands for the kernel generator $G_{\phi}$ with noise schedule as Table 2 describes. The results show that the “k-diff” mode outperforms the “Standard” mode in terms of PSNR and SSIM. The results indicate that increasing the depth of the kernel generator network leads to a more accurate kernel estimation.
Figure 4: Ablaion study on the network for PSF estimation. This figure compares the performance of the “Standard” mode against the “Diffusion” mode architecture with a varying number of hidden layers ($n=1$ to $n=5$).
Noise degradation:
To evaluate the robustness of our method to additive noise, we test it across noise levels $\sigma\in{0.01,0.02,0.03,0.05}$ using the degradation model $\mathbf{y}=\mathbf{x}\circledast\mathbf{k}+\sigma\mathbf{n},~\mathbf{n}\sim\mathcal{N}(0,\mathbf{I}).$ The image is normalized to $[0,1]$. As shown in Figure 5, our method remains stable for noise levels up to $\sigma=0.03$, achieving PSNR above $28.0$ and SSIM above $0.73$. The recovered PSFs also remain visually accurate up to $\sigma=0.03$.
Figure 5: Performance comparison under different addictive noise on the blurred image.
5 Evaluation
In this section, we evaluated DeblurSDI on datasets described in Section 4.1 comparing to other methods. The joint estimation of the image denoiser and the blur kernel often collapses to trivial solutions, such as Dirac kernels or reproducing the blurred image itself, especially when the chosen kernel size is incompatible with the image content. Larger kernels are harder to be recovered accurately, while smaller kernels may fail to capture long-range motion. For this reason, SelfDeblur [31] carefully selects the kernel size for each image. In contrast, our method exhibits much greater robustness. As shown in Figure 6, we evaluate ten different kernel sizes from 15 to 33, and compare the performance of several approaches. Our method not only achieves consistently superior performance across all kernel sizes but also demonstrates remarkable stability.
Optical Aberration Correction:
Based on the simulated six optical aberrations from 3.1, all the clean images are blurred by each one of them to obtain the corresponding blurred images. Then the proposed DeblurSDI and all the compared methods are applied and evaluated on the blurred images. The results are shown in Figure 7 and the quantitative results are shown in Table 4. It can be seen that the proposed DeblurSDI outperforms all the compared methods in all the optical aberration situations.
Figure 7: The reconstruction results of DeblurSDI on the optical aberration situations.
Table 4: Quantitative results of optical aberration correction performance (PSNR/SSIM) on four datasets.
Phase-Only[28] FFT-ReLU Deblur[2] SelfDeblur[31] FastDiffusionEM[21] DeblurSDI (Ours)
Levin[22] 15.52/0.3722 19.57/0.5655 18.13/0.4706 18.68/0.5085 28.36/0.8598
Kohler[19] 27.37/0.7889 29.89/0.8358 20.76/0.5409 19.83/0.5242 32.07/0.9061
FFHQ[17] 26.31/0.7703 23.21/0.6942 19.65/0.5591 17.90/0.4508 33.00/0.9343
Motion Deblur:
Figure 8 shows deblurring results of different methods on the FFHQ dataset [17]. For each estimated image, the recovered kernel is displayed at the top-left corner, and three zoomed-in regions highlight fine details. As observed, FastDiffusionEM [21] performs the worst. Despite being pre-trained on the FFHQ dataset, its performance remains poor: the estimated kernels degenerate into trivial point- or line-like structures. Without reliable kernel estimation, even a strong image prior cannot yield satisfactory deblurring. In contrast, other four optimization-based methods produce visibly superior results.
Figure 8: Deblurring results on the FFHQ dataset [17].
Figure 9: Deblurring results on the Cho dataset [8].
Figure 10: Deblurring results on the Kohler dataset [19].
Figure 11: Deblurring results on the Levin dataset [22].
PhaseOnly [28] and FFTReLU Deblur [2] provide good results in some cases, but lack of robustness when facing various blur kernels. SelfDeblur [31] shows the most promising performance out of previous methods, yet still suffers from poor generalizability, reconstruction shiftting and color distortion. The proposed DeblurSDI method consistently outperforms all compared methods, achieving significant improvements in the long-standing shiftting and robustness issues of blind deblurring techniques. Moreover, DeblurSDI is also able to recover accurate blur kernels, which gurantees the performance of the method. More results of other datasets are shown in Figure 9, 10, and 11.
Table 5: Quantitative results of blind motion deblurring performance (PSNR/SSIM).
Phase-Only[28] FFT-ReLU Deblur[2] SelfDeblur[31] FastDiffusionEM[21] DeblurSDI (Ours)
Levin[22] 20.68/0.6061 15.56/0.3845 25.06/0.7301 16.55/0.4005 31.85/0.7911
Cho[8] 19.89/0.6746 18.73/0.6546 20.37/0.6844 15.39/0.4687 28.73/0.8859
Kohler[19] 28.23/0.8092 25.33/0.7140 21.97/0.5995 18.85/0.4813 29.17/0.7653
FFHQ[17] 25.80/0.7904 21.71/0.6579 19.82/0.5563 15.59/0.3592 33.90/0.9064
We compared our DeblurSDI method with several other blind deblurring approaches, including Phase-Only [28], FFT-ReLU Deblur [2], SelfDeblur [31], and FastDiffusionEM [21] and calculated Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM) on each dataset. The quantitative results are summarized in Table 5. This demonstrates the effectiveness and generalizability of our self-diffusion approach in recovering sharp images and accurate blur kernels.
6 Discussion and Conclusion
Through solving blind inverse problem to recover the image under optical aberrations and motion blur, we investigate a novel and robust solution for image deblurring. The plausible performance and robustness of our proposed method are brought in by the noise schedule, which is a key component of the self-diffusion framework. We astutely observed this characteristic and boldly extended it to solving extremely unstable joint optimization inverse problems. This greatly solves the problem of unstable solutions or collased in trivial solutions. Of course, because we simultaneously constrained both the image and the kernel with neural networks and added a hierarchical noise schedule, our method requires a longer runtime compared to others. However, this time investment is worthwhile because we significantly improved the realism of the recovered image across the continuous spectrum, especially high-frequency details.
This paper presents a novel self-diffusion-based approach for optical aberration correction and motion deblurring, which we call DeblurSDI. Our method leverages the self-diffusion principle to recover accurate blur kernels and sharp images in a single framework. Experimental results on four benchmark datasets as well as six simulated optical aberrations show that DeblurSDI consistently outperforms other blind deblurring methods on various datasets, demonstrating its effectiveness and generalizability in recovering sharp images and accurate blur kernels.